# Scaling Laws for Pre-training Agents and World Models

### v1

@@ Abstract
The performance of embodied agents has been shown to improve by increasing model parameters, dataset size, and compute. This has been demonstrated in domains from robotics to video games, when generative learning objectives on offline datasets (pre-training) are used to model an agent’s behavior (imitation learning) or their environment (world modeling). This paper characterizes the role of scale in these tasks more precisely. Going beyond the simple intuition that ‘bigger is better’, we show that the same types of power laws found in language modeling (e.g. between loss and optimal model size), also arise in world modeling and imitation learning. However, the coefficients of these laws are heavily influenced by the tokenizer, task & architecture – this has important implications on the optimal sizing of models and data.
@@ 1 Introduction
The role of scale in pre-training is until now best understood in the context of large language models (LLMs). Following the observation that the empirical relationship between loss and key scaling quantities can be accurately described by power laws [Kaplan et al., 2020], ensuing work studied the precise trade-off between model and dataset size [Hoffmann et al., 2022], as well as considerations about inference compute [Sardana and Frankle, 2023], repeated training data [Muennighoff et al., 2024], parameter counting [Pearce and Song, 2024], and more (Section 2).
In comparison, less is understood about scaling in embodied AI. Recent high-impact works suggest increasing model and dataset size can lead to ever more capable agents for two pre-training objectives; behavior cloning (BC) [Reed et al., 2022, Baker et al., 2022, Brohan et al., 2023] and world modeling [Hafner et al., 2020, Hu et al., 2023, Yang et al., 2023, Bruce et al., 2024]. Such works typically demonstrate the benefit of scale through ablations over only a few model sizes, shown in terms of downstream agent performance, confirming the intuition that ‘bigger is better’ (Sartor and Thompson [2024] provide an aggregated analysis). However, this leaves a large gap to the precise understanding of scale in LLMs, where for a given increase in compute, models can be sized optimally, and their expected performance accurately predicted.
This paper helps close this gap. Similar to the initial study of scale in LLMs, we focus on the effect of scaling on a generative pre-training loss (rather than on downstream agent performance, or reward- or representation-centric objectives), in the infinite data regime, on a fixed offline dataset. Under this setting, we train families of transformers on next-token prediction tasks using architectures popular in both world modeling and BC tasks. This leads to several contributions, summarized in Figure 1.
@@
Organization. Section 2 provides detailed related work, contrasting the current understanding of scaling in embodied AI with other domains. Section 3 introduces details for our main experiments, including the architectures considered and details of scaling laws analyses. Section 4.1 & 4.2 present our main results in world modeling and BC respectively. Section 5 presents insights behind our main results, including a set of tiny-scale language experiments mimicking aspects of our main experiments. Section 6 discusses our findings and notes limitations in our work.
a) WM-Token-256, $N_{\text{optimal}}\propto C^{0.49}$                 b) WM-Token-540, $N_{\text{optimal}}\propto C^{0.62}$
c) BC-Token-540, $N_{\text{optimal}}\propto C^{0.32}$                   d) BC-CNN, $N_{\text{optimal}}\propto C^{0.66}$
Figure 1: This paper observes that scaling laws, as originally found in LLMs, also emerge in the tasks of world modeling and BC, when studying the pre-training loss on large datasets of human behavior. a, b) For world modeling, the power law coefficient determining optimal model size is affected by the compression rate of the tokenizer. c) In BC with tokenized image observations (BC-Token), small models need a large FLOPs budget to saturate, making these scaling laws less clear cut. d) However, moving to a single continuous embedding per observation remedies this (BC-CNN), producing prototypical scaling laws and a more balanced optimal model size coefficient.
2 Related work
Scaling laws origin. The term scaling laws is used throughout the engineering and physical sciences to denote power law relationships between two quantities, e.g. duration of a volcanic eruption and the probability of it continuing [Cannavò and Nunnari, 2016]. The name derives from the scale-invariant [^1] For two variables $x$ $y$, the power law $y=ax^{b}$ is invariant to scaling $x$ by a constant $c$.
@@ 2 Related work
property of power laws. While early work suggested that power laws could be good empirical descriptors of important variables in deep learning [Hestness et al., 2017, Rosenfeld et al., 2019], it was Kaplan et al. [2020] who provided a comprehensive study of power laws in transformer LLMs, and popularized the usage of scaling laws in this context.
@@
Scaling laws in LLMs. As the real-world value of LLMs was understood, scaling in LLMs became a high-priority research topic. Hoffmann et al. [2022] conducted a precise analysis into the trade-off of model and dataset size, finding they should be increased in equal proportions. This conflicted with Kaplan et al.’s suggestion that model size should be prioritized – an incorrect conclusion that Pearce and Song [2024] showed largely arose from counting only non-embedding parameters.
@@
Many other aspects of LLM scaling analyses are beginning to be refined. Su et al. [2024] revisited the methodology used to find scaling coefficients. Hägele et al. [2024] found that multiple independent cosine schedules could be reproduced more efficiently through a constant learning rate with multiple short decays, or stochastic weight averaging. Pearce and Song [2024] & Porian et al. [2024] found that well-tuned constant learning rates were sufficient to recover certain coefficients. Bi et al. [2024] study the effect of various hyperparameters on scaling. Muennighoff et al. [2024] looked at repeated epochs, finding up to four epochs produce negligible departures from the infinite data regime. Sardana and Frankle [2023] factored in inference compute to the definition of what is compute-optimal. Isik et al. [2024] study the link between pre-training loss and downstream performance. A further line of research aims to explain why power laws are such a good descriptor of empirical deep learning [Hutter, 2021, Maloney et al., 2022, Bahri et al., 2024].
@@
Scaling in embodied AI. Compared to LLMs, our understanding of scale in embodied settings is less advanced. Early successes in competitive games showed that running reinforcement learning (RL) at scale could surpass human performance, e.g. [Silver et al., 2017, Berner et al., 2019]. In self-play RL, power laws were observed between certain quantities by Neumann and Gros [2022]. Meanwhile, Hilton et al. [2023] noted that, in general, reward signals do not follow power laws, and defined a transformation of reward (intrinsic performance) that created self-consistent scaling laws.
@@
Inspired by the effectiveness of scaling in LLMs, embodied AI research has recently begun to explore the effectiveness of generative pre-training objectives on offline datasets, when executed at scale. This includes behavior cloning objectives in video games [Baker et al., 2022, Raad et al., 2024], robotics [Brohan et al., 2022, 2023, Padalkar et al., 2023, Bousmalis et al., 2023], or multiple domains [Reed et al., 2022], as well as world modeling objectives [Hu et al., 2023, Yang et al., 2023, Bruce et al., 2024]. In these studies, the benefit of scale is generally shown through increasing model size on a specific downstream task of interest (e.g. measured by completion rate) – an aggregated survey is provided by Sartor and Thompson [2024].
Whilst such studies provide strong evidence that scale can be effective in embodied domains, the complexity introduced by downstream task evaluation makes quantifying the effects, and asking more nuanced questions very challenging (e.g. how to trade-off model and dataset size). To address this, we argue that a better starting point for studying scale in these generative pre-training approaches, is to study the effect of scale on pre-training loss. Whilst downstream task performance may be of ultimate interest, LLM research evidences the utility of focusing on this clean intermediate signal, which is more straightforward to analyze. Few prior works have taken such an approach. In Appendix A we contrast our BC efforts with Tuyls et al. [2023] who provide a valuable initial study on the relationship between compute and pre-training loss for BC. We also discuss scaling analyses of image and video modeling, which are related to world modeling.
@@ 3.1 Tasks
Given this data format, we consider two tasks. World modeling (WM) [Ha and Schmidhuber, 2018] predicts future observations from previous observations and actions. This allows an agent to explicitly understand how its environment works, which can be used for planning, or dyna-style RL [Sutton, 2018]. Behavior cloning (BC) predicts the future actions that the dataset’s demonstrators take [Bakker et al., 1996]. This creates a policy that can be directly used to act in the environment, either as-is or following further fine-tuning. Concretely, these two tasks require modeling the following quantities,
@@
This work focuses on generative pre-training aiming to model this full conditional probability distribution. We leave a study of scaling laws for alternative objectives, e.g., explicitly targeting representation learning [Nair et al., 2022] or reward-centric models [Hafner et al., 2020], to future work.
@@ 3.2 Architectures
All experiments revolve around GPT-2 style causal transformers [Radford et al., 2019] as the core of the model. However we consider two different methods for inputting image observations, summarized in Figure 2. Section 3.4 details how we measure the model size of each.
@@
This tokenized architecture is widely used both in world modeling [Micheli et al., 2022] and BC tasks [Bousmalis et al., 2023]. Gato [Reed et al., 2022] used a similar design but with continuous patches rather than discrete tokens. Our implementation tests both a ‘small’ (28M parameters, $d_{z}=256$) and ‘large’ (150M parameters, $d_{z}=540$) VQGAN – further details in Appendix B.
@@
This produces an architecture similar to Baker et al. [2022] (VPT additionally used a transformer-XL and a refined hierarchical action space). Our implementation uses an Impala-style [Espeholt et al., 2018] CNN with 0.6M parameters for embedding image observations.
@@ 3.3 Datasets
    Dataset size. Repeated training on the same data alters the effect of scaling. Therefore, datasets should be large enough that all model sizes use a low number of training epochs.
@@
    Dataset diversity. Both the behavior and environment must contain enough richness and variety that pre-training loss does not saturate across the model sizes tested.
@@
Our work instead focuses on a dataset of human behavior collected in a video game named Bleeding Edge. This is a fast-paced 4 vs 4 multiplayer game, with a range of characters, abilities and maps. Game play is highly complex due to the cooperative and competitive dynamics. Success requires selecting high-level strategies (e.g. choosing which map regions to fight for), as well as fine-grained reactive control during combat. Figure 9 shows example sequences from our dataset.
Supported by the game’s developer Ninja Theory, we compiled a dataset of 8.6 years of anonymized game play, containing both image observations and controller actions. We refer to this as the 7 map dataset. We also use a subset of this for some experiments, of around 1.1 years from a single map, which we name the Sky Garden dataset. Appendix B.3 provides further technical details.
@@ 3.4 Scaling analysis methodology
We are interested in studying the relationship between several quantities defined below.
  • Model size $N$, the total number of trainable parameters (ignoring VQGAN parameters for WM-Token & BC-Token, but including the fixed-size CNN for BC-CNN). Embedding parameters are included in the count following Pearce and Song [2024].
@@
  • Compute $C$, the number of floating point operations (FLOPs) used during training. The common approximation of $C=6ND$ [Kaplan et al., 2020] is used.
  • Loss $L$, the standard classification cross-entropy loss (all targets are discretized). We assume training loss is an accurate proxy for test loss (Appendix B.3.1 analyzes further).
@@
$\hat{N}_{\text{optimal}}(C)$ $=a_{0}C^{-a}\;\;\;\;\;\;\hat{D}_{\text{optimal}}(C)=b_{0}C^{-b},$ (6)
with fitted constants $a_{0},a,b_{0},b$. We consider two methods to fit these relationships, introduced by Hoffmann et al. [2022]. Their Method 1, which we term Frontier fit, classifies efficient models as those falling on the efficient frontier (see Figure 1). Coefficients can then be estimated straightforwardly through a line of best fit on a plot of FLOPs vs parameters or data for these efficient models.
@@
In these situations, we resort to Method 3 of Hoffmann et al. [2022], which we term Parametric fit. This fits the coefficients $\alpha,\beta,N_{c},D_{c},E$ to a parametric loss form,
@@
to the empirical training curves. In our implementation, we use SciPy’s curve_fit function. We then find $a=\beta/(\alpha+\beta),b=\alpha/(\alpha+\beta)$. This makes a very strong assumption about the training curves, but allows coefficients to be estimated at a smaller compute budget.
For loss prediction we use the form recommended by Pearce and Song [2024],
@@
Training details. While early scaling studies conducted sweeps over multiple cosine decays of differing lengths [Kaplan et al., 2020, Hoffmann et al., 2022], follow up work found this redundant [Pearce and Song, 2024, Hägele et al., 2024, Porian et al., 2024]. We follow the approach of using a constant learning rate per model, so each requires only one training run. We aim to train models until they have passed their compute efficient FLOPs budget. We only modify the parameters of the transformer, following the configurations documented in Appendix B.

### v2

# library.bib

@@ Abstract
The performance of embodied agents has been shown to improve by increasing model parameters, dataset size, and compute. This has been demonstrated in domains from robotics to video games, when generative learning objectives on offline datasets (pre-training) are used to model an agent’s behavior (imitation learning) or their environment (world modeling). This paper characterizes the role of scale in these tasks more precisely. Going beyond the simple intuition that ‘bigger is better’, we show that the same types of power laws found in language modeling also arise in world modeling and imitation learning (e.g. between loss and optimal model size). However, the coefficients of these laws are heavily influenced by the tokenizer, task & architecture – this has important implications on the optimal sizing of models and data.
@@ 1 Introduction
The role of scale in pre-training is until now best understood in the context of large language models (LLMs). Following the observation that the empirical relationship between loss and key scaling quantities can be accurately described by power laws (kaplan2020scaling), ensuing work studied the precise trade-off between model and dataset size (hoffmann2022training), as well as considerations about inference compute (sardana2023beyondchinchilla), repeated training data (muennighoff2024repeatscaling), parameter counting (pearce2024reconciling), and more (Section 2.1).
In comparison, less is understood about scaling in embodied AI. Recent high-impact works show increasing model and dataset size can lead to ever more capable agents for two pre-training objectives; behavior cloning (BC) (reed2022gato; baker2022vpt; brohan2023rt2) and world modeling (WM) (hafner2020mastering; hu2023gaia; yang2023unisim; bruce2024genie). Such works typically demonstrate the benefit of scale through ablations over a few model sizes, shown in terms of downstream agent performance, confirming the intuition that ‘bigger is better’ (sartor2024neuralscaleembodied provide an aggregated analysis). However, this leaves a large gap to the precise understanding of scale in LLMs, where for a given increase in compute, models can be sized optimally, and their performance accurately predicted.
This paper helps close this gap. Similar to the study of scale in LLMs, we focus on the effect of scaling on a generative pre-training loss (rather than on downstream agent performance, or reward- or representation-centric objectives), in the infinite data regime, on a fixed offline dataset. Under this setting, we train families of transformers on next-token prediction tasks using architectures popular in both world modeling and BC tasks. This leads to several contributions, summarized in Figure 1.
(a) WM-Token-256
[…]
(b) WM-Token-540
[…]
(c) BC-Token-540
[…]
(d) BC-CNN
[…]
Figure 1: This paper observes that scaling laws, as originally found in LLMs, also emerge in the tasks of world modeling and BC, when studying the pre-training loss on large datasets of human behavior. (a) (b) For world modeling, the power law coefficient determining optimal model size is affected by the compression rate of the tokenizer. (c) In BC with tokenized image observations (BC-Token), small models need a large FLOPs budget to saturate, making these scaling laws less clear cut. (d) However, moving to a single continuous embedding per observation remedies this (BC-CNN), producing prototypical scaling laws and a more balanced optimal model size coefficient.
@@
Organization. Section 2.1 provides detailed related work, contrasting the current understanding of scaling in embodied AI with other domains, and justifying pre-training loss as a proxy for online reward. Section 3 introduces details for our main experiments, including the architectures & datasets considered, and details of scaling laws analyses. Section 4 presents our main results in world modeling and BC. Section 5 presents insights behind our main results, including a set of tiny-scale language experiments mimicking aspects of our main experiments. Section 6 discusses our findings and notes limitations.
2 Background
This section introduces related work, and outlines arguments and evidence supporting using pre-training loss to study scaling in embodied AI.
2.1 Related work
Scaling laws origin. The term scaling laws is used throughout the engineering and physical sciences to denote power law relationships between two quantities, e.g. duration of a volcanic eruption and the probability of it continuing (cannavo2016volcanic). The name derives from the scale-invariant [^1] For two variables $x$ $y$, the power law $y=ax^{b}$ is invariant to scaling $x$ by a constant $c$.
@@ 2.1 Related work
property of power laws. While early work suggested that power laws could be good empirical descriptors of pre-training loss in deep learning (hestness2017deep; rosenfeld2019constructive), it was kaplan2020scaling who provided a comprehensive study of power laws in transformer LLMs, and popularized the usage of scaling laws in this context.
@@
Scaling laws in LLMs. As the real-world value of LLMs was understood, scaling in LLMs became a high-priority research topic. hoffmann2022training conducted a precise analysis into the trade-off of model and dataset size, finding they should be increased in equal proportions. This conflicted with the suggestion that model size should be prioritized kaplan2020scaling – an incorrect conclusion that pearce2024reconciling showed largely arose from counting only non-embedding parameters.
@@
Many other aspects of LLM scaling analyses are beginning to be refined. su2024unraveling revisited the methodology used to find scaling coefficients. hagele2024scaling found that multiple independent cosine schedules could be reproduced more efficiently through a constant learning rate with multiple short decays, or stochastic weight averaging. pearce2024reconciling & porian2024resolving found that well-tuned constant learning rates were sufficient to recover certain coefficients. bi2024deepseek study the effect of various hyperparameters on scaling. muennighoff2024repeatscaling looked at repeated epochs, finding up to four epochs produce negligible departures from the infinite data regime. sardana2023beyondchinchilla factored in inference compute to the definition of what is compute-optimal. isik2024downstream study the link between pre-training loss and downstream performance. A further line of research aims to explain why power laws are such a good descriptor of empirical deep learning (hutter2021learning; maloney2022solvable; bahri2024explaining).
Scaling laws in image and video generation. Scaling laws have also been observed in auto-regressive modeling of video and images henighan2020scaling; tian2024visual. henighan2020scaling found the optimal trade off between model and dataset size to match their reported LLM coefficient ($N_{\text{optimal}}\propto C^{0.7}$) and was not affected by tokenizer. Our experiments offer different findings in the domain of world modeling – using updated methodologies to measure this trade-off (hoffmann2022training), we find it is affected by the tokenizer.
@@
Scaling in embodied AI. Compared to LLMs, an understanding of scale in embodied settings is less advanced. Early successes in competitive games showed that running reinforcement learning (RL) at scale could surpass human performance (silver2017mastering; berner2019dota). In self-play RL, power laws were observed between certain quantities by neumann2022scaling. Meanwhile, 2023singelagentscaling noted that, in general, reward signals do not follow power laws, and defined a transformation of reward (intrinsic performance) that created self-consistent scaling laws.
@@
Inspired by the effectiveness of scaling in LLMs, embodied AI research has recently begun to explore the effectiveness of generative pre-training objectives on offline datasets, when executed at scale. This includes behavior cloning objectives in video games (baker2022vpt; raad2024scaling), robotics (brohan2022rt1; brohan2023rt2; padalkar2023openx; bousmalis2023robocat), or multiple domains (reed2022gato), as well as world modeling objectives (hu2023gaia; yang2023unisim; bruce2024genie). In these studies, the benefit of scale is generally shown through increasing model size on a specific downstream task of interest (e.g. measured by completion rate) – an aggregated survey is provided by sartor2024neuralscaleembodied.
tuyls2023scalingimitation offer a valuable initial investigation into scaling laws for BC. They fit power laws to both BC pre-training loss and online return, when scaling width of single-layer LSTM models on datasets generated by fixed high-reward policies. We extend this line of investigation by studying transformer models trained on datasets of human behavior, discovering effects of architecture choices on scaling coefficients. In addition we study scaling laws in world models for the first time.
NetHack                     Bank Heist                     Breakout                   Space Invaders
Figure 2: Our meta-analysis of tuyls2023scalingimitation evidences that pre-training loss is strongly correlated with reward in BC tasks when in the infinite data regime.
2.2 Pre-training loss as a proxy for performance
A major difference between scaling research in LLMs and embodied AI is that LLM research uses pre-training loss as the main variable of interest, while embodied AI has focused on downstream online task performance. This handicaps embodied AI scaling research – measuring online performance for a single model checkpoint in complex environments like robotics or modern video games is expensive in time and hardware, requiring multiple repeated runs to allow statistically significant comparisons. Furthermore, models may first require a period of fine-tuning before evaluation. By contrast, pre-training loss is available for free at any point of a model’s training.
We believe embodied AI’s focus stems from reports that validation loss is only weakly correlated with online performance hussenot2021hyperparameter; li2024simpler. However, such observations have been made with fixed-sized training datasets and held out validation sets, where effects of overfitting may be slightly beneficial. In contrast, scaling law studies are conducted in an infinite data regime, where datapoints are not trained on more than once, making train and test losses equivalent, and overfitting effects not applicable.
To evidence that pre-training loss can be a good proxy for online return in the infinite data regime, we conducted a meta-analysis of (tuyls2023scalingimitation), who were able to roll out a large number of checkpoints for two reasons. 1) They used simple lightweight environments (Atari & NetHack). 2) Their demonstration policy was high-skill removing any need for fine-tuning. Figure 2 plots online environment return vs. pre-training loss for several environments (computed by tabulating pairs of points from Figure 6 & 10 in (tuyls2023scalingimitation)). The correlation coefficient for all environments is stronger that -0.94. Figure 3 shares evidence from our experiments that pre-training loss is well correlated with the video-generation quality of world models – providing correlation coefficients around 0.8. Further details in Appendix D.
[…]
Figure 3: Our experiments suggest pre-training loss is a good proxy for world model quality. Further details in Appendix D.
More intuitively, improving a next-token prediction loss in BC and WM requires models to ‘know more’ about behaviors and the environment, creating more useful pre-trained checkpoints for specialization to downstream tasks. In BC, better predicting the next action in a dataset of human behavior requires understanding the objectives human’s are trying to complete, alternative social behaviors they might choose to perform, as well as making in-context inferences about the skill level and mental state of individuals. In WM, decreasing next-token prediction loss might follow a curriculum, first requiring a model to capture basic shapes and colors, then textures and physics, followed by rare object interactions, and finally even complex stochastic elements in the environment such as other intelligent agents.
@@ 3.1 Tasks
Given this data format, we consider two tasks. World modeling (WM) (ha2018world) predicts future observations from previous observations and actions. This allows an agent to explicitly understand how its environment works, which can be used for planning, or dyna-style RL (sutton2018rltextbook). Behavior cloning (BC) predicts the future actions that the dataset’s demonstrators take (bakker1996robot). This creates a policy that can be directly used to act in the environment, either as-is or following further fine-tuning. Concretely, these two tasks require modeling the following quantities,
@@
This work focuses on generative pre-training aiming to model this full conditional probability distribution. We leave a study of scaling laws for alternative objectives, e.g., explicitly targeting representation learning (nair2022r3m) or reward-centric models (hafner2020mastering), to future work.
@@ 3.2 Architectures
All experiments revolve around GPT-2 style causal transformers (radford2019gpt2) as the core of the model. However we consider two different methods for inputting image observations, summarized in Figure 4. Section 3.4 details how we measure the model size of each.
@@
This tokenized architecture is widely used both in world modeling (micheli2022iris) and BC tasks (bousmalis2023robocat). Gato (reed2022gato) used a similar design but with continuous patches rather than discrete tokens. Our implementation tests both a ‘small’ (28M parameters, $d_{z}=256$) and ‘large’ (150M parameters, $d_{z}=540$) VQGAN – further details in Appendix A.
@@
This produces an architecture similar to baker2022vpt (VPT additionally used a transformer-XL and a refined hierarchical action space). Our implementation uses an Impala-style (espeholt2018impala) CNN with 0.6M parameters for embedding image observations.
@@ 3.3 Datasets
    Dataset size. Repeated training on the same data alters the effect of scaling – datasets should be large enough that training is done in the infinite data regime.
@@
    Dataset diversity. The behavior and environment must contain enough richness and variety that pre-training loss does not saturate across model sizes tested.
@@
Our work primarily focuses on a dataset of human behavior collected in a video game Bleeding Edge. This is a fast-paced 4 vs 4 multiplayer game, with a range of characters, abilities and maps. Game play is highly complex due to the cooperative and competitive dynamics. Success requires selecting high-level strategies (e.g. choosing which map regions to fight for), as well as fine-grained reactive control during combat. Figure 5 shows example sequences from our dataset.
Supported by the game’s developer Ninja Theory, we compiled a dataset of 8.6 years of anonymized game play, containing both image observations and controller actions. We refer to this as the 7 map dataset. We also use a subset of this for some experiments, of around 1.1 years from a single map, which we name the Sky Garden dataset. Appendix A.3 provides further details.
As a secondary dataset we use RT-1 (brohan2022rt1), comprising 14 days of human’s operating a robotic arm doing a range of manipulation tasks such as ‘pick banana from white bowl’. Using this smaller dataset allows us both to verify that conclusions on the large-scale video games dataset hold in a real-world robotics domain, and also allows us to run several small scale ablations in WM. Appendix C provides further details about the dataset and tokenizers used.
@@ 3.4 Scaling analysis methodology
We study the relationship between several quantities defined below.
  • Model size $N$, the total number of trainable parameters (ignoring VQGAN parameters for WM-Token & BC-Token, but including the fixed-size CNN for BC-CNN). Embedding parameters are included in the count following pearce2024reconciling.
@@
  • Compute $C$, the number of floating point operations (FLOPs) used during training. The common approximation of $C=6ND$ (kaplan2020scaling) is used.
  • Loss $L$, the standard classification cross-entropy loss (all targets are discretized). We assume training loss is an accurate proxy for test loss (Appendix A.3.1 analyzes further).
@@
$\hat{N}_{\text{optimal}}(C)$ $=a_{0}C^{a}\;\;\;\;\;\;\hat{D}_{\text{optimal}}(C)=b_{0}C^{b},$ (6)
with fitted constants $a_{0},a,b_{0},b$ [^2] Note that by subscribing to $C=6ND$ we find $a=1-b$; $N\propto C^{a}\implies C/D\propto C^{a}\implies D\propto C^{1-a}$. Hence, at times we only describe relationships in terms of $N\propto C^{a}$, with $N\propto D^{1-a}$ implied. We consider two methods to fit these relationships, introduced by hoffmann2022training. Their Method 1, which we term Frontier fit, classifies efficient models as those falling on the efficient frontier (see Figure 1). Coefficients can then be estimated straightforwardly through a line of best fit on a plot of FLOPs vs parameters or data for these efficient models.
@@
In these situations, we resort to Method 3 of hoffmann2022training, which we term Parametric fit. This fits the coefficients $\alpha,\beta,N_{c},D_{c},E$ to a parametric loss form,
@@
to the empirical training curves. In our implementation, we use SciPy’s curve_fit function. We find $a=\beta/(\alpha+\beta),b=\alpha/(\alpha+\beta)$. This makes a very strong assumption about the training curves, but allows coefficients to be estimated at a smaller compute budget.
For loss prediction we use the form recommended by pearce2024reconciling,
@@
Training details. While early scaling studies conducted sweeps over multiple cosine decays of differing lengths (kaplan2020scaling; hoffmann2022training), follow up work found this redundant (pearce2024reconciling; hagele2024scaling; porian2024resolving). We follow the approach of using a constant learning rate per model, so each requires only one training run. We aim to train models until they have passed their compute efficient FLOPs budget. We only modify the parameters of the transformer, following the configurations documented in Appendix A.